OpenStudy (anonymous):

Let N = 999 999 999 999 999 999. How many N's are there in N^2 ???

OpenStudy (anonymous):

N

OpenStudy (anonymous):

there is N amount of N's

OpenStudy (anonymous):

If u say 1 u will be in trouble....

OpenStudy (anonymous):

Consider N = 9 N = (10 -1) => N^2 = (10-1)^2 = 10^2 +1 -2*10*1= 101 - 20 = 81 That is 101 has 1 zero and 20 has 1 zero No 9's in N^2 Consider N =99 N = (100 -1) => N^2 = (100-1)^2 = 100^2 +1 -2*100*1= 10001 - 200 = 19801 That is 10001 has 3 zero and 20 has 2 zero Number of 9's in N^2 = 3 zeros - 2zeros = 1 9's replicate it for N = 999 999 999 999 999 999 18 => 9's are there in N Therefore 17 9's will be there in N^2

OpenStudy (anonymous):

so u want the amount 9's that will b predicted by that pattern?

OpenStudy (anonymous):

U didn't ask for 9's, u asked for N's....

OpenStudy (anonymous):

im thinking 1

OpenStudy (anonymous):

I already said if he said that, I would just conclude, well u know...

OpenStudy (anonymous):

Sorry guys its how many 9 's are there in N^2???its a typo error..

OpenStudy (anonymous):

I have concluded from this and your other questions that you are not serious.

OpenStudy (anonymous):

Actually, I have reached some other conclusions as well, the rules require me to be polite though.

OpenStudy (anonymous):

I just posted the question which I am solving or trying to solve.

OpenStudy (anonymous):

You posted the wrong question and waited for many posts to correct it.

OpenStudy (anonymous):

I apologies for that...

OpenStudy (anonymous):

sorry its 17 9's, one 8, one 1 and the rest are zeros